Fifty percent of midsize manufacturers planned to have management representative visit Canada and Mexico to take advantage of opportunities created by the North American Free Trade Agreements. An export-import group in Toronto , Canada , has invited 20 US midsize manufacturers to participate in the conference to explore trade opportunities. What are the variance and standard deviation of manufacturers sending representatives?
Answer by math_tutor2020(3817)


Answers:
Variance = 5 (exact)
Standard Deviation = 2.236068 (approximate)


Explanation:
This is a binomial distribution problem.
n = 20 = sample size
p = 50% = 0.50 = probability a company sends a representative

The variance of this binomial distribution is
v = variance
v = n*p*(1-p)
v = 20*0.50*(1-0.50)
v = 5

The standard deviation is the square root of the variance
sigma = standard deviation
sigma = sqrt(v)
sigma = sqrt(5)
sigma = 2.236068
This is approximate to 6 decimal places. Round it however your teacher instructs.